Determine the cardinality of each of the sets, A, B, and C defined below, and prove the cardinality of any set that you claim is countably infinite.


A is the set of negative odd integers


B is the set of positive integers less than 1000


C is the set of positive rational numbers with numerator equal to 1



1
Expert's answer
2021-11-15T07:38:12-0500

Let us determine the cardinality of each of the sets, "A, B," and "C" defined below, and prove the cardinality of any set that we claim is countably infinite.


Recall that the set "X" is called countably infinite if it has the same cardinality as the set "\N=\{1,2,3,\ldots\}" of positive integer numbers, that is there is a bijection "f:\N\to X."


Let "A" be the set of negative odd integers, that is "A=\{1-2k:k\in\N\}." Let us show that the map "f:\N\to A,\ f(n)=1-2n," is bijection. If "f(a)=f(b)," then "1-2a=1-2b." It follows that "2a=2b," and hence "a=b." We conclude that the map "f" is injection. For any "1-2n\in A" we have that "f(n)=1-2n," and thus the map is surjection. Consequently, "f:\N\to A,\ f(n)=1-2n,"

is bijection, and we conclude that the set "A" is countably infinite.


Let "B" is the set of positive integers less than 1000. It follows that the set "A" contains "999" elements, and hence "|B|=999."


Let "C"  is the set of positive rational numbers with numerator equal to 1, that is "C=\{\frac{1}{n}:n\in\N\}."

Let us show that the map "f:\N\to C,\ f(n)=\frac{1}n," is bijection. If "f(a)=f(b)," then "\frac{1}a=\frac{1}b." It follows that "a=b," and hence the map "f" is injection. For any "\frac{1}n\in C" we have that "f(n)=\frac{1}n," and thus the map is surjection. Consequently, "f:\N\to C" is bijection, and we conclude that the set "C" is countably infinite.